OpenStudy (anonymous):

The equation y=-1777x=27,153 can be used too predict the number y of gun deaths in the United Sates x years after 2000, that is x=0 corresponds to 2000, x=2 corresponds to 2002, x=4 corresponds to 2004, and so on. Preict the number of gun death in 200 and 2007. Inwhat year will the number of gun death be 11,160. 2004= 2007= what year will it be 11,160 Please help

OpenStudy (anonymous):

there is something off about that equation....-1777 + 27153? but even that seems off maybe no negative sign?

OpenStudy (anonymous):

but with given equation x = 4 in 2004 so 4*-1777 = -7108 if that should be +27153 - 7108 = 20045 deaths in 2004 14714 in 2007 and 11160 = 27153 - 1777*x.......-27153 from both sides 11160 - 27153 = -1117*x -15993 = -1777*x......divide both sides by -1117 -15993/-1777 = x x= 9.......so 2009